OpenStudy (anonymous):

to solve2x^2-5x=3 by completing the square, I divided each side of the equation by 2, and then added the square of 5/2 to each side of the equation. am i right?

OpenStudy (lgbasallote):

no. you add both sides by the square of 5/4

OpenStudy (lgbasallote):

do you know why?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

first you divide by 2 \[x^2 - \frac 52 x = \frac 32\] then you divide 5/2 and get the square \[x^2 - \frac 52 + (\frac{\frac 52}{2})^2 = \frac 32 + (\frac{\frac 52}{2})^2\] \[x^2 - \frac 52 + (\frac{5}{2 \times 2})^2 = \frac 32 + (\frac{5}{2 \times 2})^2\] \[x^2 - \frac 32 - (\frac 54)^2 = \frac 32 + (\frac 54)^2\] do you see it now?

OpenStudy (lgbasallote):

that last line should be x^2 - 5/2 not 3/2

OpenStudy (lgbasallote):

i need to go now so i hope you understood that :D

OpenStudy (anonymous):

i get it now. . . kinda